OpenStudy (anonymous):

An airplane originally at rest on a runway accelerates uniformly at 6.0 meters per second2 for 12 seconds. During this 12-second intervai, the airplane travels a distance of approximately

OpenStudy (anonymous):

432 m

OpenStudy (anonymous):

Suvat again, so use the equation s=u^2 + 1/2at^2 u = 0 here so its just 0.5*6*(12)^2 which is 432, just to clarify if that was unclear how he got to that.

OpenStudy (anonymous):

a = 6m/s^2 initial velocity = vi = 0 t = 12s distance = s = (vi)(t) + 0.5(a)(t)^2 = (0)(12) + 0.5(6)(12)^2 = 0 + 432 = 432 m